# Extracting work from coherence in a two-mode Bose-Einstein condensate

### v1

@@ III.3 Improved protocol using squeezing
The boson number distribution of a GCS is Poissonian, and hence the mean and variance are always equal. For sufficiently large mean boson number, Eq. (24), the work output becomes dominated by $W_{\mathrm{class}}$. The classical work output can be suppressed much more effectively by squeezing the initial coherent state. Squeezing of a two-mode BEC is possible using multiple techniques [79, 35], including time evolution with a one-axis twisting Hamiltonian [80, 81, 33] and parametric downconversion via spin-changing collisions [82, 83, 84]. A squeezed coherent state is [85, 86, 87],
@@
Figure 3: (a) Work extracted from an initial squeezed coherent state, divided into quantum $\beta W_{\mathrm{quant}}$ (solid curves) and classical $\beta W_{\mathrm{class}}$ (dotted curves) contributions. Colours indicate different values of $\beta g$. The work output is dominated by $W_{\mathrm{quant}}$ for $\langle\hat{n}\rangle\lesssim n_{S}$, with $n_{S}$ given by Eq. (32) and marked by vertical dashed lines. Lighter-coloured thicker lines give the large $\langle\hat{n}\rangle$ result for $W_{\mathrm{class}}$, Eq. (30), demonstrating subextensive scaling of $W_{\mathrm{class}}$. Grey dashed line gives the large $\langle\hat{n}\rangle$ result for $W_{\mathrm{quant}}$, Eq. (31). (b) Optimised squeezing parameters $e^{-2|\zeta|}$ (solid curves, left axis) and $\cos\theta$ (dotted curves, right axis) for the squeezed coherent state (see text for details). For $\langle\hat{n}\rangle\leq 2(3\beta g)^{-3/2}$ the squeezed variance can be matched to the thermal variance with appropriate choice of $e^{-2|\zeta|}$. For larger $\langle\hat{n}\rangle$ matching is no longer possible, see Eq. (29); minimising the squeezed variance then gives $e^{-2|\zeta|}=(4\langle\hat{n}\rangle)^{-1/3}$ (gray line), independent of $\beta g$. This results in a discontinuity in the slope of $e^{-2|\zeta|}$ and $W_{\mathrm{class}}$ at $\langle\hat{n}\rangle=2(3\beta g)^{-3/2}$ (marked by vertical arrows for $\beta g=10^{-5}$).
@@
$\operatorname{min}\langle\delta\hat{n}^{2}\rangle=\left(\frac{27}{32}\right)^{1/3}\langle\hat{n}\rangle^{2/3}$ (28)
@@
$W_{\mathrm{class}}\approx g\left(\frac{27}{32}\right)^{1/3}\langle\hat{n}\rangle^{2/3},$ (30)
@@
$W_{\mathrm{quant}}\approx\frac{1}{2\beta}\left[1+\frac{1}{3}\ln\left(\frac{27\langle\hat{n}\rangle^{2}}{32}\right)\right],$ (31)
@@ IV Discussion and conclusion
We have shown how work can be extracted from coherence in a two-mode BEC. Work almost entirely from coherence can be extracted from a GCS with precisely tuned mean boson number. The protocol can be substantially improved using squeezing, enabling work extraction predominantly from coherence over a much broader range of parameters. Due to the many-body nature of the system, the work from coherence can equivalently be interpreted as work from entanglement. A challenging step in our protocol is the controlled transfer of heat to the BEC alongside particle exchange between the two modes to reach equilibrium. A possible heat source is a separate atomic gas in analogy with sympathetic cooling [100]. Particle exchange could occur via tunnelling for two trapped modes [101, 102] or via spin-changing collisions with the reservoir for two spin states [103, 104].
The Hamiltonian (7) also describes a large, fully-connected spin chain [105] and hence could be realised in other atomic systems [35], including atomic ensembles coupled to light [106, 107], Rydberg arrays [108, 109] and trapped ions [110, 111]. The Hamiltonian also resembles that of photons in a non-linear medium [88], with the important difference that we allow the free-photon spectrum to be negative, $\Delta<0$. To realise this, photon number would need to be conserved [112, 113].
We have limited our analysis to Glauber coherent and squeezed coherent states. Improved operation may be possible using self-phase modulation of the squeezed state [114] or by using non-Gaussian states [115] at the expense of more involved state generation. We have also limited our analysis to initial states with zero von Neumann entropy; relaxing this provides an interesting area for future investigation. Extracting work from coherence can enhance power output compared to classical systems [6]. Our results provide a starting point to explore this in a two-mode BEC, and may result in a many-body quantum advantage when incorporated into an engine cycle.
@@ Appendix A Choosing squeezing parameters
We choose $\alpha$ and $\zeta$ for the initial squeezed coherent state as follows. The first three cumulants of $p_{n}^{S}$ are evaluated from the generating function $K(s)=\ln\sum_{n}p_{n}^{S}e^{sn}$, which can be written in closed form using Mehler’s formula [116]
@@
$\begin{split}\langle\hat{n}\rangle=&\sinh^{2}|\zeta|+|\alpha|^{2},\\ \langle\delta\hat{n}^{2}\rangle=&2\cosh^{2}|\zeta|\sinh^{2}|\zeta|+|\alpha|^{2}\left(\cosh 2|\zeta|-\sinh 2|\zeta|\cos\theta\right),\\ \langle\delta n^{3}\rangle=&\frac{1}{2}\sinh 2|\zeta|\sinh 4|\zeta|\\ &+\frac{3|\alpha|^{2}}{2}\left(\cosh 4|\zeta|-\frac{1}{3}-\sinh 4|\zeta|\cos\theta\right),\end{split}$ (37)
@@ Appendix C Minimum variance of a squeezed state
with $x_{1}$ corresponding to a local maximum of Eq. (38) and $x_{2}>x_{1}$ a local minimum. Substituting $x<x_{1}$ into the expression for $\langle\hat{n}\rangle$ (Eq. (37)) would give a negative value for $|\alpha|^{2}$, hence $x\geq x_{1}$. Therefore $x_{2}$ gives the global minimum of $\langle\delta\hat{n}^{2}\rangle$ for valid values of $x$. Substituting $x_{2}$ into Eq. (38) gives Eq. (28) for large $\langle\hat{n}\rangle$.

### v3

@@ III.3 Improved protocol using squeezing
The boson number distribution of a GCS is Poissonian, and hence the mean and variance are always equal. For sufficiently large mean boson number, Eq. (27), the work output becomes dominated by $W_{\mathrm{class}}$. The classical work output can be suppressed much more effectively by using a squeezed initial state. Squeezing of a two-mode BEC is possible using multiple techniques [81, 35], including time evolution with a one-axis twisting Hamiltonian [82, 83, 33] and parametric downconversion via spin-changing collisions [84, 85, 86]. A displaced squeezed state is [87, 88, 89],
@@
Figure 3: (a) Work extracted from an initial squeezed state, divided into quantum $\beta W_{\mathrm{quant}}$ (solid curves) and classical $\beta W_{\mathrm{class}}$ (dotted curves) contributions. Colours indicate different values of $\beta g$. The work output is dominated by $W_{\mathrm{quant}}$ for $\langle\hat{n}\rangle\lesssim n_{S}$, with $n_{S}$ given by Eq. (35) and marked by vertical dashed lines. Lighter-coloured thicker lines give the large $\langle\hat{n}\rangle$ result for $W_{\mathrm{class}}$, Eq. (33), demonstrating subextensive scaling of $W_{\mathrm{class}}$. Grey dashed line gives the large $\langle\hat{n}\rangle$ result for $W_{\mathrm{quant}}$, Eq. (34). (b) Optimised squeezing parameters $e^{-2|\zeta|}$ (solid curves, left axis) and $\cos\theta$ (dotted curves, right axis) for the squeezed state (see text for details). For $\langle\hat{n}\rangle\leq 2(3\beta g)^{-3/2}$ the squeezed variance can be matched to the thermal variance with appropriate choice of $e^{-2|\zeta|}$. For larger $\langle\hat{n}\rangle$ matching is no longer possible, see Eq. (32); minimising the squeezed variance then gives $e^{-2|\zeta|}=(4\langle\hat{n}\rangle)^{-1/3}$ (gray line), independent of $\beta g$. This results in a discontinuity in the slope of $e^{-2|\zeta|}$ and $W_{\mathrm{class}}$ at $\langle\hat{n}\rangle=2(3\beta g)^{-3/2}$ (marked by vertical arrows for $\beta g=10^{-5}$).
@@
$\operatorname{min}\langle\delta\hat{n}^{2}\rangle\approx\left(\frac{27}{32}\right)^{1/3}\langle\hat{n}\rangle^{2/3}$ (31)
@@
$W_{\mathrm{class}}\sim g\left(\frac{27}{32}\right)^{1/3}\langle\hat{n}\rangle^{2/3},$ (33)
@@
$W_{\mathrm{quant}}\approx\frac{1}{2\beta}+\frac{1}{2\beta}\ln\left[2\pi\left(\frac{27}{32}\right)^{1/3}\langle\hat{n}\rangle^{2/3}\right],$ (34)
@@ IV Discussion and conclusion
We have shown how work can be extracted from coherence in a two-mode BEC. Work almost entirely from coherence can be extracted from a GCS with precisely tuned mean boson number. The protocol can be substantially improved using squeezing, enabling work extraction predominantly from coherence over a much broader range of parameters. Due to the many-body nature of the system, the work from coherence can equivalently be interpreted as work from entanglement.
The coherent steps in our protocol can be realised in current experiments with two-mode BECs, which could be spatial or spin modes [46]. For two spatial modes the detuning can be controlled by varying trap depth [47] and populations can be coherently controlled using Josephson oscillations [47, 59]. For two spin states the detuning can be controlled using Zeeman fields [48, 49] and populations can be controlled using a Rabi pulse [60, 61]. The effect of changing $\Delta$ in the protocol could also be achieved by manipulating $g$ using a Feshbach resonance [72, 73, 74]. Squeezing of a two-mode BEC can be realised using a variety of techniques [81, 35, 82, 83, 33, 84, 85, 86]. Our protocol requires $\beta g\ll 1$, which is satisfied in most realisations of dilute gaseous BECs [52, 53].
The experimentally challenging part of the protocol is the thermalisation of the system with a reservoir (step 3 in Fig. 1). This could potentially be achieved by immersing the system in a thermal quantum gas of another species, in analogy with sympathetic cooling [102]. An optical tweezer trap would allow the system to be moved in and out of the reservoir as needed. Particle exchange between the two modes during thermalization could occur via tunnelling for two spatial modes [103, 104] or via spin-changing collisions with the reservoir for two spin states [105, 106, 107, 108]. We have assumed decoherence only occurs during the thermalising step of our protocol. Additional decoherence will likely reduce the work output by a term that scales linearly with the irreversible entropy production [75].
The Hamiltonian Eq. (7) also describes a large, fully-connected spin chain [109] and hence could be realised in other atomic systems [35], including atomic ensembles coupled to light [110, 111], Rydberg arrays [112, 113] and trapped ions [114, 115]. The Hamiltonian also resembles that of photons in a non-linear medium [90], with the important difference that we allow the free-photon spectrum to be negative, $\Delta<0$. To realise this, photon number would need to be conserved [116, 117].
We have limited our analysis to Glauber coherent and squeezed states. Improved operation may be possible using self-phase modulation of the squeezed state [118] or by using non-Gaussian states [119] at the expense of more involved state generation. We have also limited our analysis to initial states with zero von Neumann entropy; relaxing this provides an interesting area for future investigation. Extracting work from coherence can enhance power output compared to classical systems [6]. Our results provide a starting point to explore this in a two-mode BEC, and may result in a many-body quantum advantage when incorporated into an engine cycle.
@@ Appendix A Choosing squeezing parameters
We choose $\alpha$ and $\zeta$ for the initial squeezed state as follows. The first three cumulants of $p_{n}^{S}$ are evaluated from the generating function $K(s)=\ln\sum_{n}p_{n}^{S}e^{sn}$, which can be written in closed form using Mehler’s formula [120]
@@
$\begin{split}\langle\hat{n}\rangle=&\sinh^{2}|\zeta|+|\alpha|^{2},\\ \langle\delta\hat{n}^{2}\rangle=&2\cosh^{2}|\zeta|\sinh^{2}|\zeta|+|\alpha|^{2}\left[\cosh\left(2|\zeta|\right)-\sinh\left(2|\zeta|\right)\cos\theta\right],\\ \langle\delta n^{3}\rangle=&\cosh\left(2|\zeta|\right)\sinh^{2}\left(2|\zeta|\right)\\ &+\frac{3|\alpha|^{2}}{2}\left[\cosh\left(4|\zeta|\right)-\frac{1}{3}-\sinh\left(4|\zeta|\right)\cos\theta\right],\end{split}$ (40)
@@ Appendix C Minimum variance of a squeezed state
with $x_{1}$ corresponding to a local maximum of Eq. (41) and $x_{2}>x_{1}$ a local minimum. Substituting $x\leq x_{1}$ into the expression for $\langle\hat{n}\rangle$ (Eq. (40)) would give a negative value for $|\alpha|^{2}$, hence $x>x_{1}$. Therefore $x_{2}$ gives the global minimum of $\langle\delta\hat{n}^{2}\rangle$ for valid values of $x$. Substituting $x_{2}$ into Eq. (41) gives Eq. (31) for large $\langle\hat{n}\rangle$.